3. A sample of consumers was surveyed regarding their visits to a computer chain store and if the store was conveniently located. The following table summarizes the results of that survey:

Visits Convenient Not Convenient Totals
Often 60 20 80
Occasional 25 35 60
Never 5 50 55
Column Totals 90 105 195
a. What is this table called
b. Are frequency of visits and convenience independent or related? Explain your answer. What is the probability of a survey respondent being convenient to a store?
c. What is the probability of a survey respondent being not convenient to a store and never visiting a store in the chain?

Answer by stanbon(57327)

A sample of consumers was surveyed regarding their visits to a computer chain store and if the store was conveniently located. The following table summarizes the results of that survey:

Visits Convenient Not Convenient Totals
Often 60 20 80
Occasional 25 35 60
Never 5 50 55
Column Totals 90 105 195
a. What is this table called: A contingency table.
-------------------------------
b. Are frequency of visits and convenience independent or related? Explain your answer.
P(often) = 80/195 ; P(convenient) = 90/195
P(often)*P(confenient) = 0.1893
P(often | convenient) = 60/90 = 2/3
Since P(oft | conven) is not equal to P(oft)*P(conven)
they are dependent factors.
--------------------------------------------------------------

What is the probability of a survey respondent being convenient to a store?
P(convenient+ = 90/195
-----------------------------
c. What is the probability of a survey respondent being not convenient to a store and never visiting a store in the chain?
P(not convenient AND never visiting) = P(con | never)*P(never)
= 5/55 * 55/195 = 5/195
===========================
Cheers,
Stan H.
================================